OpenStudy (anonymous):

Holly has a rectangular garden that measures 12 m wide by 14 m long. She wants to increase the area to 255 m2 by increasing the width and length by the same amount. What will be the dimensions of the new garden?

OpenStudy (anonymous):

I hate doing these, geometry is not my strong suit in math ._. Have @kiwi8965 help you.

OpenStudy (anonymous):

the new dimensions of the garden will be (12+x)m wide and (14+x)m long Use the area formula A=lw (12+x)(14+x)=255m^2 x^2+26x-87=0 (x+29)(x-3)=0 x cannot be negative because she is adding to the dimensions so x=3. the new dimensions will be 15m wide x 17m long

OpenStudy (anonymous):

best response thanks

OpenStudy (anonymous):

np!